# Estimation of Out-of-Sample Sharpe Ratio for High Dimensional Portfolio Optimization

### v1

@@ Abstract
Portfolio optimization aims at constructing a realistic portfolio with significant out-of-sample performance, which is typically measured by the out-of-sample Sharpe ratio. However, due to in-sample optimism, it is inappropriate to use the in-sample estimated covariance to evaluate the out-of-sample Sharpe, especially in the high dimensional settings. In this paper, we propose a novel method to estimate the out-of-sample Sharpe ratio using only in-sample data, based on random matrix theory. Furthermore, portfolio managers can use the estimated out-of-sample Sharpe as a criterion to decide the best tuning for constructing their portfolios. Specifically, we consider the classical framework of Markowits mean-variance portfolio optimization with known mean vector and the high dimensional regime of $p/n\to c\in(0,\infty)$, where $p$ is the portfolio dimension and $n$ is the number of samples or time points. We propose to correct the sample covariance by a regularization matrix and provide a consistent estimator of its Sharpe ratio. The new estimator works well under either of three conditions: (1) bounded covariance spectrum, (2) arbitrary number of diverging spikes when $c<1$, and (3) fixed number of diverging spikes when $c\geq 1$. We can also extend the results to construct global minimum variance portfolio and correct out-of-sample efficient frontier. We demonstrate the effectiveness of our approach through comprehensive simulations and real data experiments. Our results highlight the potential of this methodology as a powerful tool for portfolio optimization in high dimensional settings.
@@
This supplementary material contains additional numerical results and all the proofs for theorems in the main text.
@@ 1 Introduction
The mean-variance portfolio optimization, first introduced by Markowitz (1952), is a fundamental cornerstone in modern portfolio allocation theory. Suppose that the excessive returns of $p$ risky assets have population excess mean vector ${\bm{\mu}}\in\mathbb{R}^{p}$ and population covariance matrix ${\bm{\Sigma}}\in\mathbb{R}^{p\times p}$. In mean-variance portfolio optimization, one solves for the allocation vector $\mathbf{w}$ that maximizes $\mathbf{w}^{\top}{\bm{\mu}}-\lambda\mathbf{w}^{\top}{\bm{\Sigma}}\mathbf{w}$, where $\lambda$ is the risk aversion. The solution $\mathbf{w}^{*}$ of mean-variance portfolio optimization is then proportional to ${\bm{\Sigma}}^{-1}{\bm{\mu}}$, that is, $\mathbf{w}^{*}\propto{\bm{\Sigma}}^{-1}{\bm{\mu}}$. It is well known that this solution achieves the largest possible Sharpe ratio.
@@
Despite the clean framework and theory of mean-variance portfolio, there are challenges in its practical applications. In practice, the population mean ${\bm{\mu}}$ and population covariance matrix ${\bm{\Sigma}}$ are unknown and have to be estimated using historical data. This leads to the phenomenon known as the “empirical optimism”, which states if we use the in-sample data both for estimation and for out-of-sample assessment, we are overly optimistic about the true out-of-sample performance of the estimated portfolio (Karoui, 2010). Thus, in order to genuinely evaluate the out-of-sample Sharpe ratio using historical data, we have to develop a valid and accurate estimator, which is the main purpose of the current paper. The estimator will help portfolio managers to achieve a portfolio with superior expected future performance.
@@
Many works in the literature have studied the estimation of ${\bm{\mu}}$ (Green et al., 2013; Fischer & Krauss, 2018; Baba Yara, 2020; Petropoulos et al., 2022; Wasserbacher & Spindler, 2022), which is often quite challenging as the signal level of financial data is weak. In comparison, the estimation of the population covariance matrix ${\bm{\Sigma}}$ is also particularly difficult not because of signal level, but due to the curse of high dimensionality. For instance, even when ${\bm{\Sigma}}$ is an identity matrix, it is widely known that the sample covariance matrix $\widehat{\bm{\Sigma}}$ is a poor estimator of ${\bm{\Sigma}}$ when the dimension of the matrix and the sample size increase to infinity proportionally—in this setting, the spectral distribution of $\widehat{\bm{\Sigma}}$ follows the MP-Law (Marchenko & Pastur, 1967), which is drastically different from that of true ${\bm{\Sigma}}$. Therefore, in high dimensional settings, the sample covariance matrix is not a reliable estimator of the population covariance matrix for mean-variance portfolio optimization (Bai et al., 2009). In this paper, we only focus on the issue caused by high dimensional covariance estimation and assumes the knowledge of ${\bm{\mu}}$ for simplicity. Note that we care about covariance estimation via the specific function of Sharpe ratio, which can be different from the majority of the literature on estimating covariance matrix itself, whose performance is typically measured using various matrix norms.
@@ 1.1 Literature review
Estimation errors in ${\bm{\mu}}$ and ${\bm{\Sigma}}$ results in a substantial deterioration in out-of-sample portfolio performance, which is typically measured by Sharpe ratio. Effort invested in improving the out-of-sample Sharpe ratio associated with Markowitz portfolio has a long history (Brown, 1976; Klein & Bawa, 1976; Jobson & Korkie, 1981; Craig MacKinlay & Pástor, 2000). More recently, Kan & Zhou (2007) introduced a three-fund rule which can improve the out-of-sample portfolio performance in the case where a risk-free asset is available. DeMiguel, Garlappi & Uppal (2009) empirically demonstrated that the “$1/N$ rule” consistently outperforms various optimally estimated portfolios across a wide range of datasets. Tu & Zhou (2011) enhanced the three-fund rule by incorporating the “$1/N$ rule”.
A second approach to enhance the out-of-sample Sharpe ratio involves integrating various regularization techniques. Bai et al. (2009); Karoui (2010) found that using in-sample data for both creating and evaluating portfolios leads to an overly optimistic assessment of their true out-of-sample performance, particularly as the number of assets grows. Recognizing the need for regularization, several studies (Jagannathan & Ma, 2003; Fan et al., 2012; Bruder et al., 2013; Ao et al., 2019) have highlighted its significance in preventing the deterioration of portfolio performance with increasing number of assets. Jagannathan & Ma (2003) empirically showed the advantage of no-short-sale constraints when estimating optimal portfolios, with Fan et al. (2012) providing theoretical support for this strategy. Additionally, Bruder et al. (2013) reviewed various regularization methods that help stabilize mean-variance allocations, including weight constraints, resampling methods, and shrinkage techniques. DeMiguel, Garlappi, Nogales & Uppal (2009) considered a generalized approach to portfolio optimization by constraining different portfolio norms. More recently, Li et al. (2022) gave a spectral correction to the sample covariance matrix to more accurately estimate the out-of-sample Sharpe ratio when $p<n$. Additionally, Kan et al. (2024) explored the asymptotic distribution of both in-sample and out-of-sample Sharpe ratios under the same condition with Li et al. (2022). Ao et al. (2019) applied a Lasso penalty in portfolio optimization and estimated the maximum Sharpe ratio considering scenarios where the ratio $p/n$ approaches a constant $c\in(0,1)$.
@@ 1.2 Our contributions
Several methodologies have been developed for estimating the out-of-sample Sharpe ratio. This raises an interesting question:
Is it possible to estimate the out-of-sample Sharpe ratio with some regularization using in-sample data? Can we then optimize the estimator over the regularization parameter to enhance the out-of-sample Sharpe ratio?
In this paper, we investigate mean-variance portfolio optimization with ridge regularization. Specifically, under the setting where the unknown population covariance matrix is estimated by the sample covariance, we develop a novel method for estimating the out-of-sample Sharpe ratio for the ridge penalized mean-variance portfolio. Our method can be applied to both the settings with and without the risk-free asset: with the risk-free asset, our approach directly proposes an estimator for the out-of-sample Sharpe ratio of the regularized mean-variance portfolio; without the risk-free asset, our approach is given in the form of estimating the efficient frontier, that is, for any given target return, instead of estimating the Sharpe ratio, we equivalently estimate the variance of the regularized mean-variance portfolio. Remarkably, the estimators given in our paper are consistent – while they can be calculated without the knowledge of the population covariance matrix, they still share the same asymptotic behavior as the portfolio optimization solution calculated with the knowledge of the population covariance matrix. Our approach optimizes portfolios using in-sample data and provides accurate performance evaluations. We summarize our contributions as follows:
@@
    We introduce an innovative in-sample method that estimates the true out-of-sample variance and thus the true out-of-sample Sharpe ratio of a portfolio, considering both scenarios with and without the risk-free asset. With no risk-free asset, we actually provide a way to accurately estimate the efficient frontier. This method provides an effective framework for evaluating portfolio performance in the future, backed by theoretical guarantees of almost surely convergence.
@@
    Additionally, our method assists in the selection of optimal regularization parameters from a predetermined set, enhancing the precision of investment decisions. By providing valid estimations for the future performances of portfolios under different regularization levels, our method allows data-driven decision making for the regularization parameters, which can be particularly useful in situations where market conditions change rapidly and allows the portfolio allocation to be updated in a timely fashion.
@@
    From a technical standpoint, our approach establishes theoretical convergence under relatively mild conditions. For instance, our assumption on the population covariance matrix ${\bm{\Sigma}}$ is less restrictive compared to the classical setting, which typically requires the bound of $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$. In our method, we allow $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ to tend towards infinity, expanding the applicability of our approach. In particular, the factor model implies a covariance structure with diverging $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ and our method still works under the factor models, as long as either $c<1$ or the number of diverging spikes is fixed.
Compared to the work with lasso regularization (Ao et al., 2019), our work differs from theirs in a few salient aspects. Firstly, we consider ridge regularization under more general and relaxed conditions allowing diverging spikes in population covariance matrix with arbitrary orders. Moreover, when factors exist, they consider holding factor portfolios besides individual stocks, although factor portfolios may not always be accessible. In contrast, we only allocate portfolio among stocks. Most importantly, they only give the estimation of the maximum value of Sharpe ratio, while we provide valid estimations for any regularization $q$, so that we can select the optimal tuning $q$. It is unclear how the Lasso tuning affects Sharpe ratio in Ao et al. (2019).
@@ 1.3 Paper organization and notations
The remainder of the paper is organized as follows. We first introduce some notations below. In Section 2 and Section 3, we give the problem settings for the regularized mean-variance portfolio optimization with and without risk free assets, and propose our methodology to estimate the Sharpe Ratio respectively. We then present simulation results in Section 4 to support our theoretical findings. Section 5 contains our demonstration about the effectiveness of our method for optimizing mean-variance portfolios using real financial return data. Finally, we conclude the paper in Section 6 and discuss potential future research. Additional experiments and proof details are provided in the supplementary material.
We employ lower case letters (e.g. $a$) for scalar values. Boldface letters are used to denote vectors and matrices (e.g. $\mathbf{a}$ and $\mathbf{A}$). For a set $I$ of indices, we define ${\bm{1}}(I)$ such that ${\bm{1}}(i)=1$ if $i\in I$, and ${\bm{1}}(i)=0$ if $i\notin I$. Hence, $\mathbf{a}(i)$ specifically refers to the $i$-th element in the vector $\mathbf{a}$. Moreover, the vector $\text{\bf 1}\in\mathbb{R}^{p}$ ($\text{\bf 1}_{n}\in\mathbb{R}^{n}$) denotes the vector with all coordinates equal to 1. For a square matrix $\mathbf{A}$, we use $\lVert\mathbf{A}\rVert_{{\mathrm{tr}}}$ and $\lVert\mathbf{A}\rVert_{{\mathop{\mathrm{op}}}}$ to denote its trace norm and operator norm, respectively, and use ${\mathrm{tr}}(\mathbf{A})$ to denote its trace. The sets of natural, real and complex numbers are denoted by $\mathbb{N}$, $\mathbb{R}$ and $\mathbb{C}$, respectively. The set of integers from $n_{1}$ to $n_{2}$ is denoted by $[n_{1}:n_{2}]=\{n_{1},\ldots,n_{2}\}$ and $[n]=[1:n]=\{1,\ldots,n\}$.
@@ 2.1 Mean-variance portfolio and Sharpe ratio
In this section, we introduce the problem setting we consider in this paper. We first review mean-variance portfolio optimization under the existence of the risk-free rate $r_{0}\geq 0$.
@@
Given $p$ risky assets with mean $\mathbf{r}\in\mathbb{R}^{p}$ and covariance matrix ${\bm{\Sigma}}\in\mathbb{R}^{p\times p}$, the mean-variance portfolio optimization is to optimize the allocation vector $\mathbf{w}$, which satisfies
@@
Here, we denote by ${\bm{\mu}}=\mathbf{r}-r_{0}\text{\bf 1}$ the excess return of the risky assets, and $\mu_{0}>0$ is the targeted excess return of the portfolio.
The motivation behind Definition 2.1 is clear. Each allocation vector $\mathbf{w}$ represents a unique portfolio, and the optimal $\mathbf{w}$ in Definition 2.1 minimizes the variance of the portfolio with an expected excess return. It is worth noting that in the mean-variance optimization, it is not required to have $\mathbf{w}^{\top}{\bm{1}}=1$, because the existence of a risk-free asset makes investing or borrowing risk-less bond possible. To be exact, we could allocate $w_{0}=1-\mathbf{w}^{\top}{\bm{1}}$ to the risk-free asset, so that the total mean excess return is indeed $(\mathbf{w}^{\top}\mathbf{r}+w_{0}r_{0})-r_{0}=\mathbf{w}^{\top}{\bm{\mu}}$.
The solution in Definition 2.1 is $\mathbf{w}^{*}\propto{\bm{\Sigma}}^{-1}{\bm{\mu}}$. Although the theory behind mean-variance portfolio optimization is elegant, in practice, a portfolio manager needs to estimate ${\bm{\Sigma}}$ in order to find out the optimal mean-variance portfolio $\mathbf{w}^{*}$. However, the estimation of ${\bm{\Sigma}}$ is very challenging due to the huge number of unknown parameters. To further investigate the effect of estimation error of ${\bm{\Sigma}}$, we model the sample data as follows:
@@
In this work, we assume the knowledge of $r_{0}$, ${\bm{\mu}}$ (or equivalently $\mathbf{r}$), and only focus on the effect of estimation error from estimating ${\bm{\Sigma}}$ using the sample covariance $\widehat{\bm{\Sigma}}=\mathbf{X}^{\top}\mathbf{X}/n$, where $\mathbf{X}$ is defined in Assumption 2.2. It has been widely known that when dimension $p$ is large with $p/n\to c>0$, that is $p$ grows with the sample size linearly, the sample covariance matrix is not a good estimator for the true covariance matrix. To mitigate the bad condition number of the sample covariance, it is natural to consider adding regularization in the portfolio construction. We add a regularization term $\mathbf{Q}$ to $\widehat{\bm{\Sigma}}$ in constructing the mean-variance portfolio. Therefore, we consider the following optimization:
@@
The optimal $\mathbf{w}^{*}$ satisfies
$\mathbf{w}^{*}\propto(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}.$
Here, $\mathbf{Q}$ is a positive definite matrix. The Sharpe ratio of $\mathbf{w}^{*}$ is defined as the mean of the excessive return of the portfolio over its standard deviation. Mathematically,
@@
for some deterministic matrix $\mathbf{A}\in\mathbb{R}^{p\times p}$. Hence, by setting $\mathbf{A}={\bm{\mu}}{\bm{\mu}}^{\top}$, it is easy to see that the Sharpe Ratio (SR) in (2.1) could be expressed as
@@
We introduce a deterministic matrix $\mathbf{A}\in\mathbb{R}^{p\times p}$ to obtain results more general than the choice of $\mathbf{A}={\bm{\mu}}{\bm{\mu}}^{\top}$. Here we have the underlying assumption that the mean vector of financial returns keeps unchanged over the time period under consideration. However, if daily excess returns have time-varying means, denoted as ${\bm{\mu}}_{i}$, where $i$ represents the trading date, and if the daily returns are independent across dates, the annual out-of-sample Sharpe ratio of the daily-balanced portfolio $\mathbf{w}_{i}^{*}\propto(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}_{i}$ can be expressed via setting $\mathbf{A}=\sum_{i}{\bm{\mu}}_{i}{\bm{\mu}}_{i}^{\top}$ over all trading dates.
To obtain the largest out-of-sample Sharpe ratio, one might select the regularization matrix $\mathbf{Q}$ by a cross-validation procedure, but this is on the one hand computationally expensive, on the other hand not straightforward to do for a time series setting in practice. Note that although for theoretical analysis, we assume data are independent from time to time, but in practice, it is not recommended to shuffle time order in the computation of out-of-sample Sharpe ratio. In this work, we hope to directly propose a method to estimate Sharpe ratio using in-sample data, which can be used as a good criterion to select the best regularization $\mathbf{Q}$ to guarantee good out-of-sample Sharpe performance of the regulated mean-variance portfolio above. In the next section, without any information of the population covariance matrix ${\bm{\Sigma}}$, we propose a statistic which allows us to apply in-sample data to effectively estimate the Sharpe ratio $SR(\mathbf{Q})$ in (2.4).

### v3

@@ Abstract
Portfolio optimization aims at constructing a realistic portfolio with significant out-of-sample performance, which is typically measured by the out-of-sample Sharpe ratio. However, due to in-sample optimism, it is inappropriate to use the in-sample estimated covariance to evaluate the out-of-sample Sharpe, especially in the high dimensional settings. In this paper, we propose a novel method to estimate the out-of-sample Sharpe ratio using only in-sample data, based on random matrix theory. Furthermore, portfolio managers can use the estimated out-of-sample Sharpe as a criterion to decide the best tuning for constructing their portfolios. Specifically, we consider the classical framework of Markowits mean-variance portfolio optimization under high dimensional regime of $p/n\to c\in(0,\infty)$, where $p$ is the portfolio dimension and $n$ is the number of samples or time points. We propose to correct the sample covariance by a regularization matrix and provide a consistent estimator of its Sharpe ratio. The new estimator works well under either of the following conditions: (1) bounded covariance spectrum, (2) arbitrary number of diverging spikes when $c<1$, and (3) fixed number of diverging spikes with weak requirement on their diverging speed when $c\geq 1$. We can also extend the results to construct global minimum variance portfolio and correct out-of-sample efficient frontier. We demonstrate the effectiveness of our approach through comprehensive simulations and real data experiments. Our results highlight the potential of this methodology as a useful tool for portfolio optimization in high dimensional settings.
@@
This supplementary material contains additional experimental results and all the proofs for theorems in the main text.
@@ 1 Introduction
The mean-variance (MV) portfolio, first introduced by Markowitz (1952), is a fundamental cornerstone in modern portfolio allocation theory. Suppose that the excessive returns of $p$ risky assets have population excess mean vector ${\bm{\mu}}\in\mathbb{R}^{p}$ and population covariance matrix ${\bm{\Sigma}}\in\mathbb{R}^{p\times p}$. In mean-variance portfolio optimization, one solves for the allocation vector $\mathbf{w}$ that maximizes $\mathbf{w}^{\top}{\bm{\mu}}-\lambda\mathbf{w}^{\top}{\bm{\Sigma}}\mathbf{w}$, where $\lambda$ is the risk aversion. The solution $\mathbf{w}^{*}$ of the MV portfolio optimization is proportional to ${\bm{\Sigma}}^{-1}{\bm{\mu}}$. It is well known that this solution achieves the largest possible Sharpe ratio.
@@
Despite the clean framework of the mean-variance portfolio, the key challenge in its practical application is that, the population mean ${\bm{\mu}}$ and population covariance matrix ${\bm{\Sigma}}$ are unknown and have to be estimated using historical data. This leads to the phenomenon known as “empirical optimism”, which states if we use the in-sample data both for estimation of ${\bm{\Sigma}},\mu$ and for out-of-sample assessment, we are overly optimistic about the true out-of-sample performance of the estimated portfolio (Karoui, 2010). For example, if we have $n$ iid $p$-dimensional return vectors $\mathbf{R}_{i}\sim N({\bm{\mu}},{\bm{\Sigma}}),i=1,\dots,n$, assuming ${\bm{\mu}}$ is known and $p/n\to c<1$, the optimized MV portfolio based on the sample covariance $\widehat{\bm{\Sigma}}=\frac{1}{n}\sum_{i=1}^{n}(\mathbf{R}_{i}-{\bm{\mu}})(\mathbf{R}_{i}-{\bm{\mu}})^{\top}$, i.e. $\mathbf{w}^{*}=C\widehat{\bm{\Sigma}}^{-1}{\bm{\mu}}$ for some $C>0$, has the in-sample Sharpe ratio of $\sqrt{{\bm{\mu}}^{\top}\widehat{\bm{\Sigma}}^{-1}{\bm{\mu}}}$ while its out-of-sample Sharpe ratio is ${\bm{\mu}}^{\top}\widehat{\bm{\Sigma}}^{-1}{\bm{\mu}}/\sqrt{{\bm{\mu}}^{\top}\widehat{\bm{\Sigma}}^{-1}{\bm{\Sigma}}\widehat{\bm{\Sigma}}^{-1}{\bm{\mu}}}$ (see (2.1) for Sharpe ratio calculation). According to our Theorem 2.7 (with $\mathbf{Q}=\text{\bf 0}$), we will see that the latter is approximately $(1-c)\sqrt{{\bm{\mu}}^{\top}\widehat{\bm{\Sigma}}^{-1}{\bm{\mu}}}$, so that the in-sample Sharpe ratio is $1/(1-c)$ times larger than the actual Sharpe ratio. When $c$ is close to $1$, the portfolio performance will be significantly exaggerated. Through this example, we see the importance of developing a valid and accurate estimator to genuinely evaluate the out-of-sample Sharpe ratio using historical data. In this paper, we will propose such an estimator to help portfolio managers to achieve the desired portfolio performance.
@@
The example above highlights the importance of addressing estimation errors in the population parameters. Although abundant literature studied the estimation of ${\bm{\mu}}$ (Green et al., 2013; Fischer & Krauss, 2018; Baba Yara, 2020; Petropoulos et al., 2022), which is challenging as the signal level of financial data is weak, we will mainly focus on the issue caused by estimating the high-dimensional population covariance matrix ${\bm{\Sigma}}$, i.e. the curse of high dimensionality. When ${\bm{\Sigma}}$ is an identity matrix, it is widely known that the sample covariance matrix $\widehat{\bm{\Sigma}}$ is a poor estimator of ${\bm{\Sigma}}$ when the dimension of the matrix and the sample size increase to infinity proportionally—in this setting, the spectral distribution of $\widehat{\bm{\Sigma}}$ follows the MP-Law (Marchenko & Pastur, 1967; Bai et al., 2009), which is drastically different from that of true ${\bm{\Sigma}}$. Therefore, in high dimensional mean-variance portfolio optimization, we would not directly plug in the unreliable sample covariance matrix. As a remedy, in this paper, we consider adding a ridge-type of penalty to the MV optimization to alleviate the big estimation error from the sample covariance. Note that we care about covariance estimation via the specific function of Sharpe ratio, which is different from the majority of the literature on estimating covariance matrix itself, whose performance is typically measured using various matrix norms.
@@ 1.1 Literature review
Estimation errors in ${\bm{\mu}}$ and ${\bm{\Sigma}}$ result in a substantial deterioration in out-of-sample portfolio performance, which is typically measured by Sharpe ratio. Effort invested in improving the out-of-sample Sharpe ratio associated with Markowitz portfolio has a long history (Brown, 1976; Klein & Bawa, 1976; Jobson & Korkie, 1981; Craig MacKinlay & Pástor, 2000). More recently, Kan & Zhou (2007) introduced a three-fund rule which can improve the out-of-sample portfolio performance in the case where a risk-free asset is available. DeMiguel, Garlappi & Uppal (2009) empirically demonstrated that the “$1/N$ rule” consistently outperforms various optimally estimated portfolios across a wide range of datasets. Tu & Zhou (2011) enhanced the three-fund rule by incorporating the “$1/N$ rule”.
A second approach to enhance the out-of-sample Sharpe ratio involves integrating various regularization techniques. Bai et al. (2009); Karoui (2010) found that using in-sample data for both creating and evaluating portfolios leads to an overly optimistic assessment of their true out-of-sample performance, particularly as the number of assets grows. Recognizing the need for regularization, several studies (Jagannathan & Ma, 2003; Fan et al., 2012; Bruder et al., 2013; Ao et al., 2019) have highlighted its significance in preventing the deterioration of portfolio performance with increasing number of assets. Jagannathan & Ma (2003) empirically showed the advantage of no-short-sale constraints when estimating optimal portfolios, with Fan et al. (2012) providing theoretical support for this strategy. Additionally, Bruder et al. (2013) reviewed various regularization methods that help stabilize mean-variance allocations, including weight constraints, resampling methods, and shrinkage techniques. DeMiguel, Garlappi, Nogales & Uppal (2009) considered a generalized approach to portfolio optimization by constraining different portfolio norms. More recently, Li et al. (2022) gave a spectral correction to the sample covariance matrix to more accurately estimate the out-of-sample Sharpe ratio when $p<n$. Additionally, Kan et al. (2024) explored the asymptotic distribution of both in-sample and out-of-sample Sharpe ratios under the same condition with Li et al. (2022). Ao et al. (2019) applied a Lasso penalty in portfolio optimization and estimated the maximal Sharpe ratio considering scenarios where the ratio $p/n$ approaches a constant $c\in(0,1)$.
@@ 1.2 Our contributions
In this paper, we investigate the MV portfolio optimization with ridge regularization. Specifically, under the setting where the unknown population covariance matrix is estimated by the sample covariance, we develop a novel method for estimating the out-of-sample Sharpe ratio for the ridge penalized MV portfolio. Our method can be applied to both the settings with and without the risk-free asset: with the risk-free asset, our approach directly proposes an estimator for the out-of-sample Sharpe ratio of the regularized MV portfolio; without the risk-free asset, our approach is given in the form of estimating the efficient frontier, that is, for any given target return, we estimate the variance of the regularized MV portfolio. Remarkably, the estimators given in our paper are consistent—while they can be calculated without the knowledge of the population covariance matrix, they still share the same asymptotic behavior as if they are calculated with the knowledge of the population covariance. We summarize our contributions as follows:
@@
    We introduce an innovative in-sample method that estimates the true out-of-sample variance and thus the true out-of-sample Sharpe ratio of a portfolio, considering both scenarios with and without the risk-free asset. With no risk-free asset, we actually provide a way to accurately estimate the efficient frontier. Our method is backed by theoretical guarantees of almost sure convergence.
@@
    Our method assists in the selection of optimal regularization parameters from a predetermined set. By providing valid estimations for the future performances of portfolios under different regularization levels, our method allows data-driven decision making for the regularization parameters, which can be particularly useful in situations where market conditions change rapidly and investment decisions must be made timely.
@@
    Our theoretical convergence results allow much relaxed assumptions on the population covariance matrix ${\bm{\Sigma}}$. Traditional assumptions in random matrix theory require $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ to be finite, while our framework allows $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$ to be unbounded with diverging eigenvalues. This makes our proposed estimators generally applicable to high-dimensional factor models with either strong or weak factors. This is achieved by rigorously handling the Stieltjes transform as $z\to 0$ and proving the interchangeability of the limits $z\to 0$ and $n\to+\infty$. Our method remains effective even with diverging $\|{\bm{\Sigma}}\|_{{\mathop{\mathrm{op}}}}$, provided that either $c<1$ or the number of diverging spikes is fixed with weak requirement on the diverging speed.
We hope to clarify that this work is neither meant to compare different portfolio optimization settings (e.g., ridge penalty versus Lasso penalty, or Value-at-risk loss versus MV objective), nor to study the optimal design of the ridge regularization (although we will comment on this in Section 2.2). Instead, we study the exact asymptotic behavior of adding the Ridge penalty to the classical mean-variance optimization so that our asymptotic results can be used for tuning the parameters in the Ridge penalty. This is in contrast to non-asymptotic study of concentration bounds, where typically only the proper rates for the tuning parameters are derived leaving the effect of their constants undetermined. Therefore, those results cannot be used to tune the tuning parameters. For example, Ao et al. (2019) applied Lasso regularization in the portfolio optimization. Our work differs from theirs in a few salient aspects. Firstly, we consider ridge regularization under more general and relaxed conditions allowing diverging spikes in the population covariance matrix with arbitrary orders. Moreover, they consider holding factor portfolios besides individual stocks. In contrast, since factor portfolios may not be always accessible, we only allocate portfolio among stocks. Most importantly, they only consider $p<n$ and give the estimation for the maximal Sharpe ratio, while we provide valid estimations for any regularization level, so that we can select the optimal tuning. It is unclear how the Lasso tuning affects Sharpe ratio in Ao et al. (2019).
@@ 1.3 Paper organization and notations
The remainder of the paper is organized as follows. We first introduce some notations below. In Section 2 and Section 3, we give the problem settings for the regularized MV portfolio optimization with and without risk free assets, and propose our methodology to estimate the Sharpe ratio respectively. We then present simulation results in Section 4 to support our theoretical findings. Section 5 contains our demonstration about the effectiveness of our method for optimizing the MV portfolio using real financial return data. Finally, we conclude the paper in Section 6 and discuss potential future research. Additional experiments and proof details are provided in the supplementary material.
[…]
@@ 2.1 Mean-variance portfolio and Sharpe ratio
In this section, we introduce the problem setting we consider in this paper. We first review the MV portfolio optimization under the existence of the risk-free rate $r_{0}\geq 0$.
@@
Given $p$ risky assets with mean $\mathbf{r}\in\mathbb{R}^{p}$ and covariance matrix ${\bm{\Sigma}}\in\mathbb{R}^{p\times p}$, the MV optimization optimizes the allocation vector $\mathbf{w}$, which satisfies
@@
Here, we denote by ${\bm{\mu}}=\mathbf{r}-r_{0}\text{\bf 1}$ the excess return of the risky assets, and $\mu_{0}>0$ is the target excess return of the portfolio.
The motivation behind Definition 2.1 is clear. Each allocation vector $\mathbf{w}$ represents a unique portfolio, and the optimal $\mathbf{w}$ in Definition 2.1 minimizes the variance of the portfolio with an expected excess return. It is worth noting that in the MV optimization, it is not required to have $\mathbf{w}^{\top}{\bm{1}}=1$, because the existence of a risk-free asset makes investing or borrowing risk-less bond possible. To be exact, we could allocate $w_{0}=1-\mathbf{w}^{\top}{\bm{1}}$ to the risk-free asset, so that the total mean excess return is indeed $(\mathbf{w}^{\top}\mathbf{r}+w_{0}r_{0})-r_{0}=\mathbf{w}^{\top}{\bm{\mu}}$.
The solution in Definition 2.1 satisfies $\mathbf{w}^{*}\propto{\bm{\Sigma}}^{-1}{\bm{\mu}}$. Here, vector $\mathbf{a}\propto\mathbf{b}$ means that $\mathbf{a}=C\mathbf{b}$ for some $C>0$. In practice, a portfolio manager needs to estimate ${\bm{\Sigma}}$ in order to find out the optimal mean-variance portfolio $\mathbf{w}^{*}$. However, the estimation of ${\bm{\Sigma}}$ is very challenging due to its huge number of unknown parameters. To further investigate the effect of estimation error of ${\bm{\Sigma}}$, we assume the following data generation process.
@@
For now, we assume the knowledge of ${\bm{\mu}}$ (or equivalently $\mathbf{r}$) 111 We first consider a known ${\bm{\mu}}$ because (1) it makes the theoretical analysis more transparent as it separates the randomness from estimating ${\bm{\mu}}$; (2) it reflects that hedge fund portfolio managers usually use many independent alternative datasets and complicated machine learning techniques to estimate ${\bm{\mu}}$ independent of the estimation of ${\bm{\Sigma}}$. However, we also consider the out-of-sample SR estimation when ${\bm{\mu}}$ is unknown and estimated by the historical sample mean; see our Remark 2.9 and Appendix G for details., and focus only on the effect of estimation error from estimating ${\bm{\Sigma}}$ using the sample covariance $\widehat{\bm{\Sigma}}=\mathbf{X}^{\top}\mathbf{X}/n$, where $\mathbf{X}$ is defined in Assumption 2.2. It has been widely known that when dimension $p$ is large with $p/n\to c>0$, that is $p$ grows with the sample size linearly, the sample covariance matrix is not a good estimator for the true covariance matrix. To mitigate the bad condition number of the sample covariance, it is natural to consider adding regularization in the portfolio construction. We add a regularization term $\mathbf{Q}$ to $\widehat{\bm{\Sigma}}$ in the MV portfolio. Therefore, we consider the following optimization:
@@
The optimal $\mathbf{w}^{*}$ satisfies $\mathbf{w}^{*}\propto(\widehat{\bm{\Sigma}}+\mathbf{Q})^{-1}{\bm{\mu}}$. Here, $\mathbf{Q}$ is a positive definite matrix. The Sharpe Ratio (SR) of $\mathbf{w}^{*}$ is defined as the mean of the excessive return of the portfolio over its standard deviation. Mathematically,
@@
for some deterministic matrix $\mathbf{A}\in\mathbb{R}^{p\times p}$. Hence, by setting $\mathbf{A}={\bm{\mu}}{\bm{\mu}}^{\top}$, it is easy to see that the Sharpe ratio in (2.1) could be expressed as
@@
We introduce a deterministic matrix $\mathbf{A}\in\mathbb{R}^{p\times p}$ to obtain results more general than $\mathbf{A}={\bm{\mu}}{\bm{\mu}}^{\top}$. Here we have the underlying assumption that the mean vector of financial returns keeps unchanged over the time period under consideration. However, if daily excess returns have time-varying means, denoted as ${\bm{\mu}}_{i}$, where $i$ represents the trading date, and if the daily returns are independent across dates, the annual out-of-sample Sharpe ratio of the daily-balanced portfolio can be expressed via setting $\mathbf{A}=\sum_{i}{\bm{\mu}}_{i}{\bm{\mu}}_{i}^{\top}$ over all trading dates.
To obtain the largest out-of-sample Sharpe ratio, one might select the regularization matrix $\mathbf{Q}$ by a cross-validation procedure, but this is on the one hand computationally expensive, on the other hand not straightforward to do for a time series setting in practice. Note that although for theoretical analysis, we assume data are independent from time to time, but in practice, it is not recommended to shuffle time order in the computation of Sharpe ratio. In this work, we hope to propose a method to directly estimate Sharpe ratio using in-sample data, which can be used as a good criterion to select the best regularization $\mathbf{Q}$ to guarantee good out-of-sample Sharpe ratio of the regulated MV portfolio.